MEASURING THE NUCLEAR HYPERFINE POPULATIONS IN THE GROUND ELECTRONIC STATE OF ATOMIC HYDROGEN USING VELOCITY-ALIGNED DOPPLER SPECTROSCOPY

K. A. Cowen, K. T. Lorenz, \underline{P. E. Fleming}, M. F. Herman, and B. Koplitz, Department of Chemistry, Tulane University, New Orleans, LA 70118.

A variation on the technique of velocity-aligned Doppler spectroscopy has been used to measure the relative populations in the F=0 and F=1 nuclear hyperfine levels in the ground electronic state of atomic hydrogen produced in the 193 nm photolysis of HBr. This technique provides an unambiguous correlation between the H atom state distributions and the spin-orbit states of the counter fragment, i.e. Br(^{2}P_{3/2}) and Br(^{2}P_{1/2}). H atom state distributions are monitored by observing the (1+1) two-color multiphoton ionization spectrum exciting through the Lyman-α transition. Effects of small Stark fields on the observed spectra will be discussed.